OpenStudy (anonymous):

Find the equation of the line that passes through the point (-3, -8) with a slope of -5. Write your answer in Slope-Intercept Form

OpenStudy (ash2326):

Bailey I showed you this yesterday also!! \[y=mx+c\] m= slope

OpenStudy (anonymous):

so, is it y=-3/8x-5 ??

OpenStudy (ash2326):

no slope m=-5

OpenStudy (ash2326):

so \[y=-5x+c\] plug in the points and find c

OpenStudy (anonymous):

y=-5x+-3/8 ?!

OpenStudy (ash2326):

tracy please don't post answer, let her try:)

OpenStudy (ash2326):

no Bailey Let me show you

OpenStudy (anonymous):

You know that X has to decrease by 5 every time Y goes up by 1, so try to find what Y is when X is equal to 0. Or you can draw a graph to find the intercept.

OpenStudy (ash2326):

we have \[y=-5x+c\] line passes through (-3, -8) x=-3 , y=-8 so \[-8=-5\times(-3) +c\] find c and tell what it is?

OpenStudy (anonymous):

-3???

OpenStudy (ash2326):

we have \[-8=15+c\] subract both sides by 15 \[-8-15=0+c\] \[c=-23\]

OpenStudy (anonymous):

y=-5x-23?????????????/

OpenStudy (ash2326):

Yeah , next time do try it first:)